# UniRVQA: A Unified Framework for Retrieval-Augmented Vision Question Answering via Self-Reflective Joint Training

### v1

@@ Abstract.
Knowledge-based Vision Question Answering (KB-VQA) systems address complex visual-grounded questions requiring external knowledge, such as web-sourced encyclopedia articles. Existing methods often use sequential and separate frameworks for the retriever and the generator with limited parametric knowledge sharing. However, since both retrieval and generation tasks require accurate understanding of contextual and external information, such separation can potentially lead to suboptimal system performance. Another key challenge is the integration of multimodal information. General-purpose multimodal pre-trained models, while adept at multimodal representation learning, struggle with fine-grained retrieval required for knowledge-intensive visual questions. Recent specialized pre-trained models mitigate the issue, but are computationally expensive. To bridge the gap, we propose a Unified Retrieval-Augmented VQA framework UniRVQA. UniRVQA adapts general multimodal pre-trained models for fine-grained knowledge-intensive tasks within a unified framework, enabling cross-task parametric knowledge sharing and the extension of existing multimodal representation learning capability. We further introduce a reflective-answering mechanism that allows the model to explicitly evaluate and refine its knowledge boundary. Additionally, we integrate late interaction into the retrieval-augmented generation joint training process to enhance fine-grained understanding of queries and documents. Our approach achieves competitive performance against state-of-the-art models, delivering a significant 4.7% improvement in answering accuracy, and brings an average 7.5% boost in base MLLMs’ VQA performance, all within a total training time of under 3 hours. The code is available at https://anonymous.4open.science/r/UniRVQA-D8C7.
@@ 1. Introduction
Knowledge-based Vision Question Answering (KB-VQA) is a task of answering challenging image-grounded questions that require the integration of external knowledge beyond commonsense, such as encyclopedic documents from the web (Vrandečić and Krötzsch, 2014; Luo et al., 2021). One of the approaches to KB-VQA involves leveraging the implicit knowledge of large pre-trained models (Yang et al., 2022; Alayrac et al., 2022; Chen et al., 2023b; Ma et al., 2024). However, these large models are less flexible in updating with the latest knowledge and often fail to mitigate their inherent factual errors (Mallen et al., 2023; Min et al., 2023). Alternatively, Retrieval-Augmented Generation (RAG) methods have emerged as a more flexible and lightweight solution for the KB-VQA task. RAG-based systems achieve knowledge-intensive generation by relying on the supporting knowledge that retrieved from external sources, providing the generator with relevant context and reducing the need for large-scale tuning.
@@
Figure 1. A comparison between the large-model-based framework, multi-stage RAG framework, and our proposed unified RAG framework (UniRVQA). The representative systems are listed in gray texts for exemplifications.
@@
Existing work on RAG-based KB-VQA often employs sequential and independent models for the retriever and answer generator (Gui et al., 2021; Lin et al., 2022, 2023). Such modular separation inherently prevents the retriever and generator from benefiting each other’s training process and mutually sharing parametric knowledge. Recent efforts  (Lin and Byrne, 2022; Siriwardhana et al., 2023) have demonstrated promise in mitigating the issue by improving the retriever’s performance through shared optimization with the generator. However, cross-task interactions and effective parametric knowledge sharing are remain constrained. These approaches retain modules separated and operate unidirectional knowledge sharing: the retriever’s parametric knowledge is refined through the generator’s training signals, without providing reciprocal feedback or sharing parametric knowledge with the generator. On the other hand, knowledge retrieval and answer generation are highly related tasks, where both of them require the precise understanding and reasoning over external knowledge documents, visual questions context, and their interrelations. We argue that a more integrated framework, facilitating mutual parametric knowledge sharing, could better exploit the interdependence between tasks and lead to improved system performance.
@@
The other fundamental challenge of KB-VQA lies in the multimodal representation learning. Earlier verbalization approaches often convert images into text (Lin et al., 2022; Gui et al., 2021; Gao et al., 2022), by image captioning and dense labeling for example. The vision-grounded task will then be reformulated as a standard textual question-answering problem. However, such transformation can result in the loss of complex visual details, thereby limiting the ability in context understanding and knowledge retrieval (Ma et al., 2024; Lin et al., 2023). With recent advancements in multimodal alignment techniques (Radford et al., 2021; Li et al., 2022), general pre-trained Multimodal Large Language Models (MLLMs), such as BLIP2 (Li et al., 2023) and InstructBLIP (Dai et al., 2023), have demonstrated their superior abilities in extracting multimodal representations and answering general visual questions. However, these pre-trained models often fail to achieve satisfactory retrieval and answering performance on knowledge-intensive bases because they are less capable of capturing fine-grained nuances. While efforts have been put into pre-training specialized models for effective knowledge-based multimodal retrieval and question answering (Chen et al., 2022; Yasunaga et al., 2022; Hu et al., 2022; Lin et al., 2023, 2024), it would be highly beneficial with promising potential to adapt existing general pre-trained multimodal models for RAG-based KB-VQA as they already memorized extensive general knowledge. Such adaptation requires much fewer data and computational resources compared to pre-trained specialized models, yet it is currently remains underexplored.
In response, this paper proposes the Unified Retrieval-Augmented Vision Question Answer (UniRVQA) framework. As demonstrated in Fig. 1-B, our framework enables a pre-trained MLLM to effectively handle all of tasks along the KB-VQA pipeline through shared parametric knowledge. Specifically, UniRVQA jointly optimizes the objectives of reflective-answering and retrieval-augmented generation. In the reflective answering branch, the model is trained to utilize its implicit knowledge repository for answering questions. A reflective-answering mechanism is proposed to enable the model to evaluate the correctness of its answers immediately after generation. During inference, this reflection mechanism determines whether external knowledge should be engaged. In the retrieval-augmented generation pathway, the process is framed as Bayesian joint probability prediction, which enables the simultaneous training of fine-grained knowledge retrieval and knowledge-dependent answer generation tasks. Additionally, we integrate late interaction mechanism (Khattab and Zaharia, 2020) into the joint training framework to further enhance fine-grained level information understanding. The experiment shows that the aforementioned tasks can complement each other, which results in improved performance in both retrieval and VQA. The main contributions of this paper are summarized as:
@@
  • We propose a novel joint training framework that enables the general MLLM to be a unified framework, which can effectively handle both tasks along the KB-VQA pipeline, including knowledge-dependent visual question answering and fine-grained knowledge retrieval.
  • We introduce a novel reflective-answering mechanism that empowers the model to assess its knowledge boundary and adaptively perform RAG during inference.
@@
  • Experiments on two public datasets show that our model outperforms state-of-the-art methods with significant improvements of 4.78% in answer accuracy and improves base MLLMs by an average of 7.54% in answering accuracy.
@@ 2.1. Knowledge-based VQA Systems
KB-VQA systems (Yang et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Yu et al., 2023; Salemi et al., 2023) solve complex knowledge-intensive visual questions. One of the approaches to KB-VQA is to leverage implicit knowledge from large language models, such as GPT-3 (Brown et al., 2020), using carefully crafted prompts. For example, KAT (Gui et al., 2021) and PICa (Yang et al., 2022) transform images into textual captions so that the visual context can be utilized as part of the prompts by GPT-3(Brown et al., 2020). However, these large models are less adaptable when updating with new knowledge and often fail to mitigate their inherent factual errors (Mallen et al., 2023; Min et al., 2023). Therefore, the retrieval-augmented generation (RAG) based paradigm has emerged as a more efficient and lightweight alternative for KB-VQA. RAG-based systems first retrieve relevant knowledge, which is then processed by the generator for answer generation. These external knowledge can either come from structured Knowledge Graphs (Li et al., 2020a; Guo et al., 2022; Speer et al., 2016) or unstructured documents such as documents on Wikipedia  (Vrandečić and Krötzsch, 2014; Lin et al., 2023).
Existing works on RAG-based KB-VQA often have independent retriever and generator models that are trained separately (Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Hu et al., 2022). For example, REVIVE (Lin et al., 2022) adopts a pre-trained CLIP (Liu et al., 2023) to extract features for retrieval and adopt multiple FiD networks  (Izacard and Grave, 2021) as the backbone of the generator. However, the independence of the generator and retriever can lead to compromised performance as their tasks are closely related. To mitigate the issue, RA-VQA (Lin and Byrne, 2022) propose to first train the retrieval network, followed by joint training of both the generator and retriever. However, cross-task interactions and parametric knowledge sharing are still constraint with the separated modules. Given the interdependence across KB-VQA tasks, we suggest that it would be beneficial to have a more integrated framework as both tasks need a fine-grained understanding of knowledge and questions. Moreover, existing work (Lin and Byrne, 2022; Yang et al., 2022; Gao et al., 2022) often verbalize images by applying image-to-text transformation (Shi et al., 2017; Anderson et al., 2018; Li et al., 2020b), which often results in the loss of critical fine-grained visual information.
@@ 3. Methodology
Before introducing the proposed framework, Section 3.1 provides a formal mathematical formulation of the KB-VQA task. To address the challenge of limited knowledge sharing between the retriever and generator, our method UniRVQA leverages a unified encoder-decoder structure built upon a base pre-trained MLLM. The encoder processes both textual and visual inputs into a unified semantic space. Section 3.3 details our novel self-reflective joint training approach, focusing on two key pathways: retrieval-augmented generation and reflective answering. Finally, we describe how answers are reasoned out, emphasizing how the reflective-answering mechanism facilitates adaptive RAG generation during inference.
@@ 3.2. Unified Multimodal Embedding
3.3. Self-Reflective Joint Training
Instead of adopting the conventional framework where the retriever and generator are separate, we adopts a unified framework, which is trained under the proposed self-reflective joint training method to optimize for retrieval and answer generation simultaneously, allowing them to complement each other during training.
@@ 3.3.1. Late-Interaction Knowledge Retrieval
During the retrieval stage, the relevance $r$ between a particular query $\mathcal{\tilde{Q}}$ and a document $\mathcal{\tilde{D}}$ is assessed using a relevance score in a late-interaction manner, following ColBERT (Khattab and Zaharia, 2020). Late-interaction retrieval is known as a fine-grained and more efficient approach, where query and document representations are independently encoded before interacting (Khattab and Zaharia, 2020). We extend the ColBERT from a BERT-based retrieval model to any MLLM-based framework that can be jointly trained with the generation task:
@@
(6) $\mathcal{L}_{R}=-\sum_{\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}}\log\frac{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D^{+}}})}{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+})+\sum_{\mathcal{\tilde{D}}^{+}\in D_{n}}\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+})}$
@@ 3.3.2. Retrieval-Augmented Generation
After the late-interaction relevance calculation, the embeddings of the positive document will be concatenated with the query embeddings and fed to the language model for answer generation. The concatenated retrieval-augmented embedding will be denoted as $y$ and the answer generation will be trained by the casual language modelling loss:
(7) $\mathcal{L}_{RAG}=-\sum_{i=1}^{l_{a}}y_{i}\log p(\hat{y}_{i}|y_{<i}),$
where $l_{a}$ is the length of ground truth answer $\mathcal{A}$. For each open-ended question, there will be a set of human responses $S$. The target answer will be randomly selected from the set. Our empirical studies show that random selection enhances the model’s robustness to document noise compared to using the answer provided in the document. Based on Eq. 2, the two loss terms can be directly summed together as the log-joint probability of retrieval and generation:
@@
1:MLLM (Encoder $\mathcal{E}$, $\mathcal{D}$, Parameters $\Phi$); Question-image pairs $q$; batch $\mathcal{B}$ with size $n$; Target Answer $\mathcal{A}$; Learning rate $\alpha$
@@
7:         $\mathcal{L}_{SR}=\mathcal{L}_{gen}=\sum_{i}^{n}\textsc{CrossEntropyLoss}(a_{i},\mathcal{A}_{i})$
@@
13:              Predict self-reflection result: $\hat{r_{i}}=\mathcal{D}(\mathcal{E}(q_{i}\oplus\hat{a_{i}}))$
@@ 3.4. Inference and Answer Select
At the inference stage, all documents will first be indexed using PLAID (Santhanam et al., 2022) for accelerated late-interaction retrieval. The self-reflection mechanism not only enables the model to evaluate its knowledge boundaries but also facilitates adaptive retrieval-augmented generation (RAG) during the UniRVQA inference stage. At this stage, the model first answers the question without referencing external documents and immediately generate self-reflection prediction to assess the correctness of its response. If the model deems its answer incorrect, the RAG process is triggered, allowing it to retrieve documents to facilitate answering. Otherwise, the self-answering result will be kept as the final answer. During the retrieval-augmented generation stage, multiple documents will be selected and generate multiple answers accordingly. Based on Eq. 1 and  2, the answer with the highest joint probability of retrieval and generation will be selected.
@@ 4.1. Experiments Setup
Datasets. We mainly evaluate the proposed method on the OK-VQA dataset (Marino et al., 2019) and conduct complementary experiments using the InfoSeek dataset (Chen et al., 2023a) to demonstrate the model’s generalizability. InfoSeek is regarded as more knowledge-dependent than earlier KB-VQA datasets, as its questions are often unanswerable without external knowledge support. Here are the details about two datasets:
(1) OK-VQA (Marino et al., 2019) dataset contains over 10k questions on MSCOCO (Lin et al., 2014) images which require external knowledge to answer. For the external knowledge base we adopt Google Search Corpus (Luo et al., 2021), which is a textual corpus, containing 166,389 passages from Google, covering all the knowledge necessary for answering questions in OK-VQA. We use the original splits of training and testing sets to ensure comparability, and use 10% of the training set for validation.
(2) InfoSeek (Chen et al., 2023a) is a newly proposed large-scale KB-VQA dataset introduced in 2023, built on the OVEN image dataset (Hu et al., 2023). Following the original paper we use Wikipedia (Vrandečić and Krötzsch, 2014) as the external knowledge base. Compared to OK-VQA, InfoSeek is more challenging as it encompasses a greater number of questions that necessitate expertise knowledge for accurate responses. Given the large size (over 1 million image-question pairs) of InfoSeek and Wikipedia corpus, we conduct our experiments on a down-sampled subset following existing works (Lin et al., 2024). Specifically, to reduce the large number of duplicate samples while maintaining its diversity, we stratified the training set by the combination of entity ID and question, and randomly sampled $\sqrt{n_{i}+1}+1$ entries from $n_{i}$ entries in each group. This sampling operation reduced the training set of Infoseek from  934k to  33k. To ensure comparability with prior works (Lin et al., 2024), we used the original validation set as the test set and randomly sample 2,000 entries from the original training set as the validation set. We also randomly sampled 100k documents from the 6M wikipedia corpus while making sure of the presence of relevant documents. Details of the data splits statistics for both datasets can be found in Table 1.
Implementation Details. We select BLIP2-Flan-T5-XL (Li et al., 2023) and InstructBLIP-Flan-T5-XL (Dai et al., 2023) as the MLLM base models to build our proposed framework. We use 1 Nvidia A100 with 80GB VRM for all experiments. We use DoRA (Liu et al., 2024) to fine-tune UniRVQA. We choose a batch size of 20 and the AdamW optimizer (Loshchilov and Hutter, 2017) with the learning rate set as 2e10-4. The scheduler modulates the learning rate throughout the training process, starting with a warmup period of 100 steps before gradually reducing the learning rate following a cosine schedule. We evaluated the models on the validation every 200 steps and the best performing checkpoints were found to be step 2800 (OK-VQA) and step 3400 (InfoSeek). The documents were truncated to a maximum length of 256 tokens. The beam size was set at 3 for answer generation. To ensure comparability and avoid randomness bias, we report our main results as the average from 3 different random seed settings.
Computation Costs. Given the best-performing checkpoints on two datasets, We report the required total training hours on 1 Nvidia A100 (80G) are on average less than 3 GPU hours. Compared to the comparable models also trained with 1 A100 and takes more than 20 GPU hours for the training progress (Lin et al., 2024, 2023), our method are less computationally expensive while achieving competitive effectiveness.
@@
(3)Pseudo Relevance Recall (PRR@K): Following previous work  (Lin and Byrne, 2022; Lin et al., 2023), we adopt pseudo-relevance labels and evaluate the retrieval performance by counting the number of questions that successfully retrieve documents with correct answers contained in top-k retrieval results.
Baseline Models. We compare our proposed framework with the latest KB-VQA systems in answer generation performance. Among them, the first group of systems smaller model with less than 3B parameters, including:
  • KAT (Gui et al., 2021) integrates implicit and explicit knowledge into a transformer-based KB-VQA system to jointly reason over both knowledge sources.
  • TRiG (Gao et al., 2022) transforms all visual information into language space on three levels, including image-level captioning, object-level dense labeling and text OCR.
  • MAVEx (Wu et al., 2022) enhances KB-VQA performance by validating generated answers through retrieving multimodal external evidence and providing multimodal explanations to support its reasoning process.
  • RA-VQA (Lin and Byrne, 2022) also transforms visual content into textual space and proposes a joint training framework for retrieval and answer generation.
  • BLIP (Li et al., 2022) is one of the leading pre-trained multimodal models, designed to perform vision-language tasks such as VQA by leveraging unified image-text semantic space.
  • FLMR (Lin et al., 2023) proposes a pre-trained framework that focuses on fine-grained alignment between multimodal inputs, enabling high-accuracy retrieval and reasoning for KB-VQA.
The second group includes larger systems that are built with large pre-trained models such as GPT-3 (175B) (Brown et al., 2020) and PaLM-E (526B):
  • PICa (Yang et al., 2022) also conducts image-text transformation and prompts the GPT-3 for implicit knowledge.
  • Prophet (Yu et al., 2023) proposes heuristics-enhanced prompting to utilize the implicit knowledge of frozen LLMs.
  • REVIVE (Lin et al., 2022) exploits object-centric regional information together with image-question query to retrieve knowledge from GPT-3 and Wikidata (Vrandečić and Krötzsch, 2014).
  • PALI (Chen et al., 2023b), Flamingo (Alayrac et al., 2022) and PaLM-E (Driess et al., 2023) represent state-of-the-art large pre-trained Multimodal models that achieve strong performance on KB-VQA datasets.
Additionally, we will compare our retrieval performance against strong retrieval-focused models designed specifically for VQA task, mainly including Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), FLMR (Lin et al., 2023) and preFLMR (Lin et al., 2024). FLMR enhances retrieval by leveraging ColBERT (Khattab and Zaharia, 2020) for fine-grained alignment, while preFLMR extends FLMR’s approach by pretraining on a substantially larger retrieval corpus to improve performance further.

### v2

# Enabling Collaborative Parametric Knowledge Calibration for Retrieval-Augmented Vision Question Answering

@@ Abstract.
Knowledge-based Vision Question Answering (KB-VQA) systems address complex visual-grounded questions with knowledge retrieved from external knowledge bases. The tasks of knowledge retrieval and answer generation tasks both necessitate precise multimodal understanding of question context and external knowledge. However, existing methods treat these two stages as separate modules with limited interaction during training, which hinders bi-directional parametric knowledge sharing, ultimately leading to suboptimal performance. To fully exploit the cross-task synergy in KB-VQA, we propose a unified retrieval-augmented VQA framework with collaborative parametric knowledge calibration. The proposed framework can effectively adapt general multimodal pre-trained models for fine-grained, knowledge-intensive tasks while enabling the retriever and generator to collaboratively enhance and share their parametric knowledge during both training and inference. To enhance fine-grained understanding of questions and external documents, we also integrate late interaction mechanism into the proposed training framework. Additionally, we introduce a reflective-answering mechanism that allows the model to explicitly evaluate and refine its knowledge boundary. Our approach achieves competitive performance against state-of-the-art models, delivering a significant 4.7% improvement in answering accuracy, and brings an average 7.5% boost in base MLLMs’ VQA performance. The code is available at https://anonymous.4open.science/r/UniRVQA-D8C7.
@@ 1. Introduction
Knowledge-based Vision Question Answering (KB-VQA) is a task of answering challenging image-grounded questions that require the integration of external knowledge beyond commonsense, such as web-sourced encyclopedic documents (Vrandečić and Krötzsch, 2014; Luo et al., 2021). Recent advances in large multimodal models demonstrate strong capabilities in leveraging the vast implicit knowledge stored in their parameters (Yang et al., 2022; Alayrac et al., 2022; Chen et al., 2023b; Ma et al., 2024). However, these large models may struggle with timely knowledge updates, factual errors and private applications (Mallen et al., 2023; Min et al., 2023). Therefore, retrieval-augmented methods have emerged as an alternative solution to the KB-VQA task, which support knowledge-intensive generation with the knowledge retrieved from external sources, providing the generator with relevant context and reducing the need for large-scale tuning.
@@
Figure 1. A comparison between the large-model-based framework, multi-stage retrieval-augmented framework, and our proposed unified framework (UniRVQA). The representative systems are listed in gray texts for exemplifications.
@@
Existing work on retrieval-augmented KB-VQA often employs sequential and independent models for the knowledge retriever and answer generator (Gui et al., 2021; Lin et al., 2022, 2023). Such modular separation inherently prevents the retriever and generator from benefiting each other’s training process and sharing parametric knowledge. Recent efforts  (Lin and Byrne, 2022; Siriwardhana et al., 2023) have explored optimizing the retriever with guidance from the generator. However, these methods still operate under a unidirectional knowledge passage: while the retriever’s parametric knowledge is refined through the generator’s training signals, no reciprocal feedback is provided to the generator. On the other hand, knowledge retrieval and answer generation both require precise understanding and reasoning over question context and external knowledge documents. Therefore, we argue that enabling bidirectional parametric knowledge sharing during training and inference could better exploit the interdependence across tasks and improve system performance.
@@
The other challenge of KB-VQA lies in the fine-grained multimodal representation learning. Earlier verbalization approaches reformulate VQA into textual QA by converting visual inputs into text via image captioning or dense labeling (Lin et al., 2022; Gui et al., 2021; Gao et al., 2022). However, such transformation often discards complex visual details, thereby limiting the ability in context understanding and downstream knowledge retrieval (Ma et al., 2024; Lin et al., 2023). Recent advancements in multimodal alignment (Radford et al., 2021; Li et al., 2022) have empowered general pre-trained Multimodal Large Language Models (MLLMs), such as BLIP2 (Li et al., 2023) and InstructBLIP (Dai et al., 2023), which excel at answering general visual questions. However, these models can struggle with knowledge-intensive tasks due to limited capacity for capturing fine-grained semantic nuances. While specialized pre-trained models have been proposed for knowledge-based retrieval and VQA (Chen et al., 2022; Yasunaga et al., 2022; Hu et al., 2022; Lin et al., 2023, 2024), it would be highly beneficial to adapt existing general pre-trained multimodal models for retrieval-augmented KB-VQA as they have already memorized extensive general knowledge. Such adaptation requires much fewer computational resources compared to pre-trained specialized models, yet it remains underexplored.
To address the challenge of insufficient parametric knowledge sharing and ineffective fine-grained multimodal representation, this paper proposes a Unified Retrieval-Augmented Vision Question Answer (UniRVQA) framework. As demonstrated in Fig.1-B, our framework is designed to calibrate the knowledge space of the retriever and the answer generator through a joint optimization scheme. The proposed setup fosters collaborative parametric knowledge learning, thereby overcoming modular isolation and enabling both components to benefit from each other’s learning process. Simultaneously, the integration of a late interaction mechanism (Khattab and Zaharia, 2020) further enhances the system’s ability to capture fine-grained semantic alignments between modalities. With an additional novel reflective-answering mechanism, the model learns not only to generate answers using retrieved knowledge, but also to assess the correctness of its own responses and internally evaluate its knowledge boundaries. The main contributions of this paper are summarized as below:
@@
  • We propose a novel KB-VQA framework with collaborative parametric knowledge calibration across tasks. The framework adapts general MLLMs to effectively handle both tasks along the KB-VQA pipeline, including fine-grained knowledge retrieval and visual question answering.
  • We introduce a reflective-answering mechanism for model to assess its knowledge boundaries, alongside incorporating the late-interaction mechanism into the unified framework to enhance fine-grained multimodal representation.
@@
  • Experiments on two public benchmarks show that our model outperforms state-of-the-art methods with significant improvements of 4.78% and improves base MLLMs by an average of 7.54% in answering accuracy.
@@ 2.1. Knowledge-based VQA Systems
KB-VQA systems (Yang et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Yu et al., 2023; Salemi et al., 2023) solve complex knowledge-intensive visual questions. One of the approaches to KB-VQA is to leverage implicit knowledge from large language models, such as GPT-3 (Brown et al., 2020), using carefully crafted prompts. For example, KAT (Gui et al., 2021) and PICa (Yang et al., 2022) transform images into textual captions so that the visual context can be utilized as part of the prompts by GPT-3(Brown et al., 2020). However, these large models are less adaptable when updating with new knowledge and often fail to mitigate their inherent factual errors (Mallen et al., 2023; Min et al., 2023). Therefore, the retrieval-augmented generation (RAG) paradigm has emerged as a more efficient and flexible alternative for KB-VQA. RAG-based systems first retrieve relevant knowledge, which is then processed by the generator for answer generation. These external knowledge can either come from structured Knowledge Graphs (Li et al., 2020a; Guo et al., 2022; Speer et al., 2016) or unstructured documents such as documents on Wikipedia  (Vrandečić and Krötzsch, 2014; Lin et al., 2023).
Existing works on retrieval-augmented KB-VQA often have independent retriever and generator models that are trained separately (Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023; Hu et al., 2022). For example, REVIVE (Lin et al., 2022) adopts a pre-trained CLIP (Liu et al., 2023) to extract features for retrieval and adopt multiple FiD networks  (Izacard and Grave, 2021) as the backbone of the generator. However, the independence of the generator and retriever can lead to compromised performance as their tasks are closely related. To mitigate the issue, RA-VQA (Lin and Byrne, 2022) propose to first train the retrieval network, followed by refining the retriever under the guidance of generator’s training signal. However, cross-task interactions and parametric knowledge sharing are still constraint with the separated modules. Therefore, the generator is blind to the retriever’s evolvement. Given the interdependence across KB-VQA tasks, we suggest that it would be beneficial to have a more integrated framework as both tasks need a fine-grained understanding of knowledge and questions. Moreover, existing work (Lin and Byrne, 2022; Yang et al., 2022; Gao et al., 2022) often verbalize images by applying image-to-text transformation (Shi et al., 2017; Anderson et al., 2018; Li et al., 2020b), which often results in the loss of critical fine-grained visual information.
@@ 3.2. Unified Multimodal Embedding
3.3. Collaborative Parametric Knowledge Calibration Training Framework
To fully utilize the interplay across retrieval and generation tasks in KB-VQA, we propose a unified framework, where the retriever and the answer generator share the parametric knowledge and mutually calibrate each other during training. Specifically, the training framework consists the simultaneous learning processes of three main tasks: late-interaction knowledge retrieval, retrieval-augmented generation and reflective answering.
@@ 3.3.1. Late-Interaction Knowledge Retrieval
During the retrieval stage, the relevance $r$ between a particular query $\mathcal{\tilde{Q}}$ and a document $\mathcal{\tilde{D}}$ is assessed using a relevance score in a late-interaction manner, following ColBERT (Khattab and Zaharia, 2020). Late-interaction retrieval is known as a fine-grained and more efficient approach, where query and document representations are independently encoded before interacting (Khattab and Zaharia, 2020). We extend ColBERT from a individual retrieval model to a flexible module compatible with any MLLM-based framework, allowing end-to-end training with the generation task.
@@
(6) $\mathcal{L}_{R}=-\sum_{\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}}\log\frac{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D^{+}}}))}{\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}))+\sum_{\mathcal{\tilde{D}}^{+}\in D_{n}}\exp(r(\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}))}$
@@ 3.3.2. Retrieval-Augmented Generation
After the late-interaction relevance calculation, the embeddings of the positive document will be concatenated with the query embeddings and fed to the language model for answer generation. The answer generation will be trained by the casual language modeling loss:
(7) $\mathcal{L}_{RAG}=-\sum_{\mathcal{\tilde{Q}},\mathcal{\tilde{D}}^{+}}\log p_{\Phi}(a|\mathcal{\tilde{Q}},\mathcal{\tilde{D^{+}}}),$
For each open-ended question, there will be a set of human responses $\mathcal{A}$. The target answer $a$ will be randomly selected from the set. Our empirical studies show that random selection enhances the model’s robustness to document noise compared to using the answer provided in the document. Based on Eq. 2, the two loss terms can be directly summed together as the log-joint probability of retrieval and generation:
@@
1:MLLM (Encoder $\mathcal{E}$, $\mathcal{D}$, Parameters $\Phi$); Question-image pairs $q$; batch $\mathcal{B}$ with size $n$; Target Answer Set $\mathcal{A}$; Learning rate $\alpha$
@@
7:         $\mathcal{L}_{SR}=\mathcal{L}_{gen}=\sum_{i}^{n}\textsc{CrossEntropyLoss}(\hat{a_{i}},\mathcal{A}_{i})$
@@
13:              Predict self-reflection result: $\hat{r_{i}}=\mathcal{D}(\mathcal{E}(q_{i},\hat{a_{i}}))$
@@ 3.4. Inference and Answer Select
At the inference stage, all documents will first be indexed using PLAID (Santhanam et al., 2022) for accelerated late-interaction retrieval. The self-reflection mechanism not only enables the model to evaluate its knowledge boundaries but also facilitates adaptive retrieval-augmented generation during the UniRVQA inference stage. At this stage, the model first answers the question without referencing external documents and immediately generate self-reflection prediction to assess the correctness of its response. If the model deems its answer incorrect, the retrieval-augmented process is triggered, allowing it to retrieve documents to facilitate answering. Otherwise, the self-answering result will be kept as the final answer. During the retrieval-augmented generation stage, multiple documents will be selected and generate multiple answers accordingly. Based on Eq. 1 and  2, the answer with the highest joint probability of retrieval and generation will be selected.
@@ 4.1. Experiments Setup
Datasets. We mainly evaluate the proposed method on the OK-VQA dataset (Marino et al., 2019) and conduct complementary experiments using the InfoSeek dataset (Chen et al., 2023a). InfoSeek is regarded as a more knowledge-dependent dataset, as its questions are often unanswerable without external knowledge support. Here are the details about two datasets:
(1) OK-VQA (Marino et al., 2019) dataset contains over 10k questions on MSCOCO (Lin et al., 2014) images which require external knowledge to answer. For the external knowledge base we adopt the Google Search Corpus (Luo et al., 2021), which is a textual corpus containing 166,389 passages from Google, covering all the knowledge necessary for answering questions in OK-VQA. We use the original splits of training and testing sets to ensure comparability, and use 10% of the training set for validation.
(2) InfoSeek (Chen et al., 2023a) is a newly proposed large-scale KB-VQA dataset introduced in 2023, built on the OVEN image dataset (Hu et al., 2023). Following the original paper we use Wikipedia (Vrandečić and Krötzsch, 2014) as the external knowledge base. Compared to OK-VQA, InfoSeek is more challenging as it encompasses a greater number of questions that necessitate expertise knowledge for accurate responses. Given the large size (over 1 million image-question pairs) of InfoSeek and Wikipedia corpus, we conduct our experiments on a down-sampled subset following existing works (Lin et al., 2024). See supplementary material for more details.
Implementation Details. We select BLIP2-Flan-T5-XL (Li et al., 2023) and InstructBLIP-Flan-T5-XL (Dai et al., 2023) as the MLLM base models to build our proposed framework. We use 1 Nvidia A100 with 80GB VRM for all experiments. We use DoRA (Liu et al., 2024) to fine-tune UniRVQA. We choose a batch size of 20 and the AdamW optimizer (Loshchilov and Hutter, 2017) with the learning rate set as 2e10-4. The scheduler modulates the learning rate throughout the training process, starting with a warmup period of 100 steps before gradually reducing the learning rate following a cosine schedule. To ensure comparability and avoid randomness bias, we report our main results as the average from 3 different random seed settings.
@@
(3) Pseudo Relevance Recall (PRR@K): Following previous work  (Lin and Byrne, 2022; Lin et al., 2023), We adopt pseudo-relevance labels and evaluate retrieval performance by counting the number of questions whose top-K retrieved documents contain the correct answers. Following prior work, we set K=5 in the main experiments.
Baseline Models. We compare our proposed framework with the latest KB-VQA systems in answer generation performance. Among them, the first group of systems smaller model with less than 3B parameters, including: KAT (Gui et al., 2021), TRiG (Gao et al., 2022), MAVEx (Wu et al., 2022), RA-VQA (Lin and Byrne, 2022), BLIP (Li et al., 2022), and FLMR (Lin et al., 2023). The second group includes larger systems that are built with large pre-trained models such as GPT-3 (175B) (Brown et al., 2020) and PaLM-E (526B): PICa (Yang et al., 2022), Prophet (Yu et al., 2023), REVIVE (Lin et al., 2022), PALI (Chen et al., 2023b), Flamingo (Alayrac et al., 2022) and PaLM-E (Driess et al., 2023). Additionally, we will compare our retrieval performance against strong retrieval-focused models designed specifically for VQA task, mainly including Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), FLMR (Lin et al., 2023) and preFLMR (Lin et al., 2024).